OpenStudy (anonymous):

A circular braided room with sides of 16 feet. When the rug is in place, how much of the old floor will show?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to compute the area of a circle whose diameter is 16 feet, so we have: \[\large A = 3.14 \times {R^2}\] and R = D/2, where D, as usual, is the diameter

OpenStudy (anonymous):

3.14*16^2?

OpenStudy (michele_laino):

3.14*8^2=...?

OpenStudy (anonymous):

200.96

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, since the area of the room is: 16*16=256 feet^2 then we have to compute this difference: \[\large 256 - 200.96 = ...?\]

OpenStudy (anonymous):

55.04!

OpenStudy (michele_laino):

ok!